A 1100kg Car Accelerates From Rest To 60.0 Km/h Over A Distance Of 45 M. find the net force acting on the car.
1
Expert's answer
2013-01-30T07:47:41-0500
m = 1100 kg
v = 60 km/h
l = 45 m

Net force F = m * a = m * v^2 / (2*l) = 1100 kg * (60 km/h)^2 / (2 * 45 m) =
3395 N

Answer: 3395 N